OpenStudy (anonymous):

What number must you add to complete the square? x^2+4x=12

OpenStudy (anonymous):

4=2^2=4?

jimthompson5910 (jim_thompson5910):

4/2 = 2 2^2 = 4 again, you are correct

OpenStudy (anonymous):

aight thxs

jimthompson5910 (jim_thompson5910):

instead of saying 4 = 2^2 make a note as to how you're going from 4 to 2